# MMD-Flagger: Leveraging Maximum Mean Discrepancy to Detect Hallucinations

### v1

@@ 3.4 Discussion
Experimental results. Table 2 lists the precision and recall scores on the benchmark datasets. MMD-Flagger (concat-Gaussian) matches the best baseline in terms of recall on the LFAN-HALL dataset (Seq-Logprob), and outperforms it (again Seq-Logprob) in terms of recall and precision on the Halomi dataset. We see these results as very promising, and observe that they are robust with respect to changes in the kernel and aggregation method.
@@
4 Discussion
Limitations. The first limitation of MMD-Flagger lies in its reliance on the source of hallucination. The method is designed to detect hallucinations by identifying discrepancies between beam search and stochastic sampling outputs, as reflected in the trajectory of MMD distances. Consequently, if the sampled sequences $Y_{\text{sto}}^{\tau}$ are highly similar to the beam search output $y_{\text{beam}}$ for low temperatures, hallucinations can go undetected. It may be the case for an irreducible proportion of hallucinations.
Another limitation is the computational cost of MMD-Flagger, as it requires repeated stochastic sampling across multiple temperature settings. This cost scales with the size of LLM and the length of the generated sequence, making the method less suitable for very large models or long outputs. The computational cost of the MMD estimate can also be high, since it is quadratic in the number of generated sequences. A practical workaround is to reduce the number of samples or temperature values used when computing the MMD trajectory. However, the stability and reliability of MMD-Flagger under such reduced settings remain an open question.
Broader Impact. Our work contributes to the development of safer and more widespread applications of LLMs. Indeed, MMD-Flagger can enhance translation-related tasks by identifying hallucinations in LLMs outputs.
@@ 5 Conclusion
In this work, we introduced MMD-Flagger, a method for hallucination detection based on the Maximum Mean Discrepancy (MMD) metric. The key idea is to compare output sequences generated by beam search and stochastic sampling across varying temperature settings. By measuring the distributional distance between these sequences using MMD, MMD-Flagger achieves detection performance on par with the strongest uncertainty-based estimation methods on two machine translation benchmarks. The method’s performance is robust with respect to changes in the output vectorization and choice of the kernel.
Future work includes extending MMD-Flagger to other generation tasks such as summarization or question-answering, using LLMs such as Qwen (Yang-2024-Qwen, ) or Llama (Grattafiori-2024llama, ) as models to inspect. Benchmark datasets are available (Maynez-etal-2020-faithfulness, ; Bao-2025-naacl, ), and the sampling strategy is the same. The challenge here is the size of the models at hand and the outputs.
Another promising research direction is to investigate the interpretability of MMD-Flagger. One possibility is to leverage recent work in feature importance for MMD-based testing sutherland2017 ; mitsuzawa-2023 . This would allow to pinpoint embedding features responsible for the detection of hallucinated content by MMD-Flagger. Once such variables are located in intermediate layers, interpretability techniques such as Layerwise Relevance Propagation (Montavon-2019, ) could be applied to analyze how these variables relate to specific encoder layers or segments of the input.
@@ Acknowledgements
This work has been supported by the French government, through the NIM-ML project (ANR-21-CE23-0005-01). We extend our sincere thanks to Prof. Motonobu Kanagawa for insightful discussions and valuable feedbacks to the conceptualization of this work. We thank Dr. Jean-Flavien Bussotti and Ms. Shilpa Sunil Vishwakarma for their careful proofreading and helpful suggestions to improve the clarity and readability.
@@
This appendix is organized in the following way: Appendix A describes relabeling procedure of annotation labels of LFAN-HALL dataset. Appendix B details the experimental setup. Appendix C contains additional results related to the hallucination detection experiments (Section 3 of the main paper). Key facts regarding the datasets used for experiments are collected in Section D. Appendix E shows qualitative results of the hallucination detection by MMD-Flagger. Comparing kernel length settings are described in Appendix F. We show the MMD-Flagger hallucination detection results using vectors from intermediate layers of models in Appendix G. Finally, we investigate the variability of the proposed hallucination detection procedure in Appendix H.
[…]
@@
Appendix A LFAN-HALL Dataset Modification and Label Correction
@@
Some annotation labels in the LFAN-HALL dataset appear to be inconsistent with the definitions given by the paper. Referring to the annotation guideline provided [27] [^1] https://github.com/deep-spin/hallucinations-in-nmt/tree/main/data, the strongly detached label is applied when “there is some content that is supported by the source sentence. However, there is a significant part of the translation that is not (it conveys different meaning to the meaning in the source sentence). Importantly, the content that is hallucinatory does not reflect a genuine effort to translate any part of the source sentence”. In Table 3, we report an example which, in the original datasets, has two labels: omission (of the machine translation error) and strongly detached (of the hallucination). This example is considered to be labeled as the omission (of the machine translation) since the machine translation correctly translated the phrase of the source text A proud celebration, yet the machine translation lacks the phrase that gives cause for celebrating. Therefore, this example is not a case of hallucination.
@@
[…]
@@
Appendix B Experimental Setup
B.1 Computational Environment and Implementation Details
Experiments are done on the following computational environment: Xeon Platinum 8268 Processor with a single GPU of Tesla V100S-PCIE $32$GB and $755$GB RAM.
B.2 Translation models
Model for LFAN-HALL. The translation model is provided by the authors [27] [^2] https://github.com/deep-spin/hallucinations-in-nmt. The model is trained by the authors using the fairseq package on the WMT18 DE-EN dataset [8]. The model is a Transformer model consisting of hidden vectors of $512$ dimensions, $6$ layers for both the encoder and decoder with $8$ attention heads. The total parameter size of the model is approximately $77$M parameters.
The translation speed is roughly $2$ seconds to sample $25$ sequences. In the experiments using vectors from the intermediate layers in Appendix G, the translation speed is significantly slower: approximately $30$ seconds for sampling $25$ sequences. This is because of the necessity of a custom codebase for extracting the vectors from the intermediate layers. Since we, regardless of our best efforts, could not find interfaces of the fairseq package to access and extract the intermediate layers during the inference mode, it is necessary to implement custom code, which causes a significant slowdown.
Model for Halomi. The translation model is the multilingual translation model [10], and the model is available on the HuggingFace Hub [^3] https://huggingface.co/facebook/nllb-200-distilled-600M. The model is based on the multilingual encoder-decoder model designed for the machine translation task [^4] The base model in the transformers package is M2M100: https://huggingface.co/docs/transformers/en/model_doc/m2m_100. The model consists of $12$ layers, and each layer generates a vector of $1,024$ dimensions.
In this case, it takes approximately $4$ seconds to sample $25$ sequences.
B.3 Implementation details
MMD-Flagger. We use an open source codebase to compute the MMD distances [37] [^5] https://github.com/Kensuke-Mitsuzawa/mmd-two-sample-test-variable-selection. The codebase is implemented using PyTorch [42]. About the computational speed, MMD-Flagger requires around $1$ second for computing MMD distances of $10$ temperature parameters when the vector aggregation mode is concat and the dimension size is $D=53,760$ and the sample size of $H_{\text{sto}}^{\tau}$ is $25$. Therefore, the required time for computing an MMD distance is around $0.1$ seconds per $\tau$ parameter.
MC-Dsim. We activate the dropout of the model for simulating the stochastic process and obtaining the translation token sequences. For calculating a distance between the token sequences using the drop-out and without, we use the METEOR metric [3] of the nltk package [7] to compute the similarity score [^6] https://www.nltk.org/api/nltk.translate.meteor_score.html.
Seq-LogProb. fairseq and transformers packages both give the log probability for a generated sequence.
@@ Appendix C Additional experimental results
C.1 Additional results for the Halomi Dataset
Table 5 represents the numbers of true-positive and false-positive per annotation label.
Table 5: Detection performance per error label on the Halomi dataset. The first row represents short names of hallucination labels that represent the detachment levels of the translation text, and the second row is the total count of the annotation label. The first column shows detection methods description of which is in Section 3.3. The value formats are “true positive / false positive.” We highlight the highest values of true positive in each annotation label.
small partial full
Count 225 291 218
MMD-Flagger (avg) 187 / 1,938 239 / 1,886 180 / 1,945
MMD-Flagger (concat) 187 / 1,989 244 / 1,932 174 / 2,002
SeqLogProb 65 / 958 142 / 881 73 / 950
MC-Dsim 109 / 836 158 / 787 139 / 806
TNG 2 / 0 2 / 0 2 / 0
Table 6 shows the detection performance for each source and target language pair. Among the $18$ language pairs analyzed, we can categorize them into three types based on the availability of training resources for the translation model: low-resource, high-resource, and zero-shot languages. Among the nine languages studied, the following three are considered low-resource languages when training the translation model: kas-Deva, mni-Beng, and yor-Latn. The source and target language pairs between Spanish and Yoruba are classified as a zero-shot learning setting that generates the translation without the training data of the corresponding source and target pair. The remaining pairs consist of high-resource languages. For detailed training information, please refer to [11].
We posited that the detection performance by MMD-Flagger may correlate with the volume of training data: low, high, and zero-shot. However, as evidenced by the results presented in Table 6, the performance of MMD-Flagger in hallucination detection does not exhibit a significant relationship with the amount of data on which the model was trained. Notably, when comparing the recall scores for MMD-Flagger, the high-resource language pairs, specifically “spa-Latn to eng-Latn” yields a lower recall score. Conversely, the recall scores for zero-shot language pairs, such as yor-Latn and spa-Latn, demonstrate relatively more favorable results.
Table 6: Detection performance per language pair on the Halomi dataset. The recall and precision scores are shown in the format of “recall (precision)”. Language pairs are written with language codes of the dataset [11]. The sub-section names “High”, “Low”, “Zero-shot” represent the training data volume for training the translation model [11]. We highlight the best recall score per language pair.
Source and Target Langs Count MMD-F(avg) MMD-F(concat) SeqLogProb TNG
High-Resource Languages
arb-Arab to eng-Latn 156 0.67 (0.17) 0.75 (0.18) 0.42 (0.16) 0.00 (0.00)
deu-Latn to eng-Latn 155 0.93 (0.13) 0.87 (0.12) 0.53 (0.13) 0.00 (0.00)
eng-Latn to arb-Arab 144 0.88 (0.06) 0.75 (0.05) 0.12 (0.02) 0.00 (0.00)
eng-Latn to deu-Latn 146 0.86 (0.12) 0.79 (0.11) 0.43 (0.10) 0.00 (0.00)
eng-Latn to rus-Cyrl 146 1.00 (0.05) 0.80 (0.04) 0.40 (0.03) 0.00 (0.00)
eng-Latn to spa-Latn 153 0.73 (0.22) 0.68 (0.21) 0.00 (0.00) 0.00 (0.00)
eng-Latn to zho-Hans 160 0.69 (0.18) 0.59 (0.14) 0.55 (0.25) 0.03 (1.00)
rus-Cyrl to eng-Latn 158 0.92 (0.09) 0.83 (0.09) 0.17 (0.03) 0.00 (0.00)
spa-Latn to eng-Latn 160 0.58 (0.22) 0.55 (0.20) 0.52 (0.27) 0.00 (0.00)
zho-Hans to eng-Latn 159 0.84 (0.21) 0.81 (0.19) 0.44 (0.22) 0.03 (1.00)
Low-Resource Languages
eng-Latn to kas-Deva 184 0.90 (0.42) 0.89 (0.41) 0.38 (0.38) 0.00 (0.00)
eng-Latn to mni-Beng 197 0.82 (0.60) 0.90 (0.63) 0.49 (0.73) 0.00 (0.00)
eng-Latn to yor-Latn 195 0.90 (0.16) 0.90 (0.16) 0.41 (0.15) 0.00 (0.00)
kas-Deva to eng-Latn 151 0.79 (0.40) 0.77 (0.39) 0.47 (0.48) 0.00 (0.00)
mni-Beng to eng-Latn 152 0.89 (0.80) 0.85 (0.79) 0.45 (0.80) 0.00 (0.00)
yor-Latn to eng-Latn 146 0.82 (0.16) 0.91 (0.16) 0.45 (0.17) 0.00 (0.00)
Zero-shot Languages
yor-Latn to spa-Latn 152 0.85 (0.51) 0.85 (0.47) 0.00 (0.00) 0.00 (0.00)
spa-Latn to yor-Latn 151 0.83 (0.35) 0.87 (0.35) 0.33 (0.30) 0.00 (0.00)

### v3

@@ 3.4 Discussion
Experimental results. Table 2 lists the precision and recall scores on the benchmark datasets. Note that the result of MMD-Flagger with concat vector representations is unavailable to FaithBench, as the resulting embedding size becomes prohibitively large and difficult to handle. MMD-Flagger (concat-Gaussian) matches the best baseline in terms of recall on the LFAN-HALL dataset (Seq-Logprob). MMD-Flagger (avg-Gaussian) outperforms baselines in terms of recall and precision on the Halomi, ConstFact, FaithBench datasets. We see these results as very promising, and observe that they are robust with respect to changes in the kernel and aggregation method.
@@
4 Limitations
The first limitation of MMD-Flagger lies in its reliance on the source of hallucination. The method is designed to detect hallucinations by identifying discrepancies between a hypothesis output and stochastic sampling outputs, as reflected in the trajectory of MMD distances. Consequently, if the sampled sequences $Y_{\text{sto}}^{\tau}$ are highly similar to the hypothesis $y_{\text{hyp}}$ for low temperatures, hallucinations can go undetected. It may be the case for an irreducible proportion of hallucinations.
Another limitation is the computational cost of MMD-Flagger, as it requires repeated stochastic sampling across multiple temperature settings. This cost scales with the size of LLM and the length of the generated sequence, making the method less suitable for very large models or long outputs. A practical workaround is to reduce the number of samples or temperature values used when computing the MMD trajectory. However, the stability and reliability of MMD-Flagger under such reduced settings remain an open question. We note that the MMD estimator is characterized by the quadratic complexity; however, the total inference time is overwhelmingly dominated by the cost of LLM output generation, thus rendering the MMD calculation a negligible component of the overall computation.
A limitation of our current framework lies in the deterministic nature of argmax decoding. The MMD-Flagger method is designed to quantify deviations from a stochastic distribution, and thus is not intended to inspect hypothesis outputs generated via argmax selection. The argmax decoding deterministically selects the highest-probability token at each step, producing rigid and often repetitive sequences, in contrast to sampling-based strategies that introduce the necessary variability for our method. Nevertheless, we acknowledge that argmax decoding is still widely adopted in practice for the purposes of reproducibility in benchmarking and evaluation (Bao et al., 2025; Zhou et al., 2025; Chuang et al., 2024).
@@ 5 Conclusion
In this work, we introduced MMD-Flagger, a method for hallucination detection based on the Maximum Mean Discrepancy (MMD) metric. The key idea is to compare output sequences generated by a default decoding strategy and stochastic sampling across varying temperature settings. By measuring the distributional distance between these sequences using MMD, MMD-Flagger achieves detection performance on par with the strongest uncertainty-based estimation methods on two machine translation benchmarks. The method’s performance is robust with respect to changes in the output vectorization and choice of the kernel.
Another promising research direction is to investigate the interpretability of MMD-Flagger. One possibility is to leverage recent work in feature importance for MMD-based testing Sutherland et al. (2017); Mitsuzawa et al. (2023). This would allow to pinpoint embedding features responsible for the detection of hallucinated content by MMD-Flagger. Once such variables are located in intermediate layers, interpretability techniques such as Layerwise Relevance Propagation (Montavon et al., 2019) could be applied to analyze how these variables relate to specific encoder layers or segments of the input.
@@ Acknowledgements
This work has been supported by the French government, through the NIMML project (ANR-21-CE23-0005-01). We extend our sincere thanks to Prof. Motonobu Kanagawa for insightful discussions and valuable feedbacks to the conceptualization of this work.
@@
This appendix is organized in the following way: Appendix A details the experimental setup. Appendix B describes relabeling procedure of annotation labels of LFAN-HALL dataset. Appendix C contains additional results related to the hallucination detection experiments (Section 3 of the main paper). Appendix H shows qualitative results of the hallucination detection by MMD-Flagger. Comparing kernel functions are described in Appendix D. We show the MMD-Flagger hallucination detection results using vectors from intermediate layers of models in Appendix E. Finally, we investigate the variability of the proposed hallucination detection procedure in Appendix F.
@@
Appendix A Configurations of experiments
A.1 Computational environment
Experiments are done on the following computational environment: Xeon Platinum 8268 Processor with a single GPU of Tesla V100S-PCIE 3232GB and 755755GB RAM.
A.2 Datasets
LFAN-HALL (Guerreiro et al., 2023b) is a hallucination detection benchmark dataset for the machine translation task from German to English. The dataset consists in $3,413$ machine translations obtained by a transformer-based encoder-decoder model relying on the fairseq package Ott et al. (2019) and beam search, together with associated reference translation and human-annotated hallucination labels. Since we observed some inaccurate annotation labels, we undertook a review of the annotated labels and made corrections. The relabeling process is detailed in Appendix B, and the corrected dataset is contained in the supplementary file code.tar.gz. The dataset defines three types of hallucination labels occupying approximately $8\%$ of all translations, and we unify these three labels as “hallucination.”
Table 3 represents the label distribution in broad categories, while Table 4 shows the label distribution following (Guerreiro et al., 2023b). According to the authors’ hallucination definition, we classify $\{$ error-repetitions, error-strong, error-full }\} into hallucination and $\{$ error named-entities, error-omission }\} into mt-error.
Label Count Ratio
no-error 2,885 0.8455
hallucination 291 0.0852
mt-error 233 0.0682
hallucination,mt-error 3 0.0008
Table 4: Label distribution of the LFAN-HALL dataset. An example is assigned two labels separated by a command separator.
Label Count Ratio
no-error 2,885 0.8455
error-omission 193 0.0565
error-full 121 0.0354
error-strong 83 0.0243
error-repetitions, error-strong 47 0.0137
error-named-entities 37 0.0108
error-repetitions 29 0.0084
error-full, error-repetitions 11 0.0032
error-named-entities, error-omission 3 0.0008
error-omission, error-strong 1 0.0002
error-omission, error-repetitions 1 0.0002
error full, error-named-entities 1 0.0002
Halomi (Dale et al., 2023) is a hallucination detection benchmark dataset consisting over $3,000$ sentences in $9$ languages. Halomi consists of beam search translations manually annotated for hallucination. The dataset defines three types of hallucinations that occupy approximately $26\%$ of the dataset, and we again unify them as “hallucination.” The label statistics of the Halomi dataset are shown in Table 5.
Table 5: Hallucination label distribution of the Halomi dataset.
Class label Count Ratio
1 No hallucination 2,131 0.74
3 Partial hallucination 291 0.10
2 Small hallucination 225 0.07
4 Full hallucination 218 0.07
ConstFact (Dreyer et al., 2023) is a dataset designed for evaluating the factual consistency of abstractive summarization models. The dataset is derived from the widely-used CNN/Daily Mail (CNN/DM) summarization dataset. Its primary purpose is to assess the trade-off between abstractiveness and factual accuracy in model-generated summaries. The dataset and its code are publicly available in the authors’ repository. [^1] https://github.com/amazon-science/abstractive-factual-tradeoff
The ConstFact dataset consists of 600 human-annotated evaluation records. Each record includes a source document and a model-generated summary. Three human annotators independently evaluated each summary, providing a label that reflects its factual consistency with the original source document. For our evaluation, we define a summary as a hallucination if, at least, one of the three human annotators marked it as “non-factual”. This operational definition allows us to categorize the model-generated summaries as either factually consistent or containing hallucinations, enabling a quantitative analysis of model performance.
FaithBench (Bao et al., 2025) is a benchmark specifically created to evaluate the factual consistency of summaries generated by modern decoder-only models. The dataset provides annotations on the “faithfulness” of model-generated summaries, categorized into four distinct labels: Questionable‘, Benign, Unwanted-Intrinsic, and Unwanted-Extrinsic. This dataset is available in the authors’ public repository. [^2] https://github.com/vectara/FaithBench
For our experiments, we constructed a custom sub-dataset from FaithBench. The original ‘FaithBench‘ dataset’s summaries were generated using a greedy search (argmax) decoding strategy. We regenerated a set of hypothesis summaries using a beam search decoding strategy with the Phi-3-mini-4k model. We randomly selected $25$ source documents from the original FaithBench dataset. Using the same prompts as the original dataset, we generated new summaries for these documents with the Phi-3-mini-4k model. Following the exact annotation scheme and guidelines outlined in the FaithBench, we then labeled the faithfulness of these $25$ newly generated summaries. Our annotation process yielded a total of $11$ summaries containing hallucinations and $14$ that are factually faithful. This new, annotated dataset is included in the supplementary ‘code.tar.gz‘ archive.
A.3 Evaluated models
LFAN-HALL. The translation model is provided by the authors (Guerreiro et al., 2023b). [^3] https://github.com/deep-spin/hallucinations-in-nmt The model is trained by the authors using the fairseq package on the WMT18 DE-EN dataset (Bojar et al., 2018). The model is a Transformer model consisting of hidden vectors of $512$ dimensions, $6$ layers for both the encoder and decoder with $8$ attention heads. The total parameter size of the model is approximately 7777M parameters. The translation speed is roughly $2$ seconds to sample $25$ sequences. In the experiments using vectors from the intermediate layers in Appendix E, the translation speed is significantly slower: approximately $30$ seconds for sampling $25$ sequences. This is because of the necessity of a custom codebase for extracting the vectors from the intermediate layers. Since we, regardless of our best efforts, could not find interfaces of the fairseq package to access and extract the intermediate layers during the inference mode, it is necessary to implement custom code, which causes a significant slowdown.
Halomi. The translation model is the multilingual translation model (Costa-Jussà et al., 2022), and the model is available on the HuggingFace Hub. [^4] https://huggingface.co/facebook/nllb-200-distilled-600M The model is based on the multilingual encoder-decoder model designed for the machine translation task. [^5] The base model in the transformers package is M2M100: https://huggingface.co/docs/transformers/en/model_doc/m2m_100 The model consists of $12$ layers, and each layer generates a vector of $1,024$ dimensions. In this case, it takes approximately $4$ seconds to sample $25$ sequences.
ConstFact. We use BART, an encoder-decoder architecture proposed by Lewis et al. (2020). Following the setup instruction by (Dreyer et al., 2023), we used the bart.large.cnn variant, a large-scale model fine-tuned on the news articles of the CNN/Daily Mail summarization dataset. The model’s parameters total $406,290,432$. The model provides a word embedding vector with a length of $1,024$ for each token. The model file is distributed through the ‘fairseq‘ repository. [^6] https://github.com/facebookresearch/fairseq/blob/main/examples/bart/README.md
Dreyer et al. (2023) introduces a custom decoding method designed to control the abstractiveness of generated summaries. Since ConstFact dataset gives annotation labels on the summaries generated with the control the abstractiveness, we followed the same configuration at generating $y_{\rm hyp}$. Our stochastic sampling of collecting $Y_{\rm sto}^{\tau}$ is not with the control the abstractiveness since authors’ proposed decoding constraints were not designed for these stochastic methods.
FaithBench. The model is Phi-3-mini-4k with 3.8B parameters. This model is available on the HuggingFace Hub. [^7] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct This model provides a word embedding vector in the length of $3,072$ for each token.
A.4 Implementation details
MMD-Flagger. We use an open source codebase to compute the MMD distances (Mitsuzawa et al., 2023). [^8] https://github.com/Kensuke-Mitsuzawa/mmd-two-sample-test-variable-selection The codebase is implemented using PyTorch (Paszke et al., 2017). About the computational speed, MMD-Flagger requires around $1$ second for computing MMD distances of $10$ temperature parameters when the vector aggregation mode is concat and the dimension size is $D=53,760$ and the sample size of $H_{\text{sto}}^{\tau}$ is $25$. Therefore, the required time for computing an MMD distance is around $0.1$ seconds per $\tau$ parameter.
MC-Dsim. We activate the dropout of the model for simulating the stochastic process and obtaining the translation token sequences. For calculating a distance between the token sequences using the drop-out and without, we use the METEOR metric (Banerjee and Lavie, 2005) of the nltk package (Bird and Loper, 2004) to compute the similarity score. [^9] https://www.nltk.org/api/nltk.translate.meteor_score.html
Seq-LogProb. fairseq and transformers packages both directly give the log probability for a generated sequence.
TNG. We use the original implementation of the algorithm provided by the authors (Raunak et al., 2021). [^10] https://github.com/GEM-benchmark/NL-Augmenter/tree/b64a8efebd842848546ec51d3402ba253be12aa7/nlaugmenter/filters/oscillatory_hallucination
Appendix B LFAN-HALL Dataset Modification and Label Correction
@@
Some annotation labels in the LFAN-HALL dataset appear to be inconsistent with the definitions given by the paper. Referring to the annotation guideline provided (Guerreiro et al., 2023b) [^11] https://github.com/deep-spin/hallucinations-in-nmt/tree/main/data, the strongly detached label is applied when there is some content that is supported by the source sentence. However, there is a significant part of the translation that is not (it conveys different meaning to the meaning in the source sentence). Importantly, the content that is hallucinatory does not reflect a genuine effort to translate any part of the source sentence.
In Table 6, we report an example which, in the original datasets, has two labels: omission (of the machine translation error) and strongly detached (of the hallucination). This example is considered to be labeled as the omission (of the machine translation) since the machine translation correctly translated the phrase of the source text A proud celebration, yet the machine translation lacks the phrase that gives cause for celebrating. Therefore, this example is not a case of hallucination.
@@
Verified $\backslash$ Orig. full full&NE full&rep NE&strong omit&rep omit&strong rep rep&strong strong
@@ Appendix C Additional experimental results
C.1 Comparison with SelfCheckGPT
We compare the hallucination detection performance of MMD-Flagger against the consistency-based method SelfCheckGPT (Manakul et al., 2023). A key difference is that SelfCheckGPT requires external models (e.g., RoBERTa, T5) for calculating its final score, whereas our MMD-Flagger approach is entirely self-contained, relying only on the internal representations of the generating model.
Setup. We evaluated the BertScore and MQAG variants of SelfCheckGPT, following the authors’ recommended protocol. For a fair comparison, we used the exact same set of $25$ stochastic samples ($\tau=1.0$) across both MMD-Flagger and SelfCheckGPT. To adapt SelfCheckGPT for the binary classification task, we set three distinct thresholds at the $25$-th, $50$-th, and $75$-th percentiles of the generated scores. Due to the high computational cost associated with generating scores for the SelfCheckGPT baseline, all reported metrics were evaluated on a carefully selected data subset; results for MMD-Flagger, MC-DSIM, and Seq-Logprob were correspondingly down-sampled for a direct comparison.
Table 8: The score formats are “Recall (Precision)”. S-GPT is SelfCheckGPT with the threshold value in the parenthesis. MC-Dsim and Seq-Logprob on ConstFact are currently unavailable.
LFAN-HALL Halomi ConstFact
MC-Dsim 0.826 (0.494) 0.612 (0.316)
Seq-Logprob 0.883 (0.481) 0.548 (0.105)
S-GPT BertScore (25) 1.000 (0.281) 1.000 (0.082) 1.000 (0.565)
S-GPT BertScore (50) 1.000 (0.317) 1.000 (0.082) 1.000 (0.575)
S-GPT BertScore (75) 0.948 (0.464) 0.903 (0.084) 0.115 (0.666)
S-GPT MQAG (25) 0.577 (0.312) 0.709 (0.118) 0.536 (0.660)
S-GPT MQAG (50) 0.242 (0.325) 0.322 (0.131) 0.014 (1.000)
S-GPT MQAG (75) 0.051 (0.350) 0.096 (0.176) 0.000 (0.000)
MMD-Flagger (avg-Gaussian) 0.808 (0.362) 0.677 (0.087) 0.855 (0.590)
MMD-Flagger (avg-Dot) 0.845 (0.357) 0.709 (0.092) 0.855 (0.590)
MMD-Flagger (concat-Gaussian) 0.856 (0.370) 0.645 (0.081) 0.884 (0.693)
MMD-Flagger (concat-Dot) 0.852 (0.375) 0.774 (0.100) 0.855 (0.686)
Results Analysis. Detection performance is summarized in Table 8, where scores are reported as Recall (Precision). While SelfCheckGPT frequently achieves a perfect recall of $1.000$ in several low-threshold settings across all three datasets (e.g., S-GPT BertScore (25)), MMD-Flagger overall achieves highly competitive Recall scores to some of SelfCheckGPT’s configurations (MQAG). The core finding is that MMD-Flagger demonstrates strong efficacy, generating comparable detection metrics without the reliance on external scoring models, thereby presenting a more self-contained solution.